OpenStudy (anonymous):

find domain of the function g(x)=1over 7-8x

OpenStudy (anonymous):

\[7-8x \neq 0 \]\[7\neq8x\]\[x \neq 7/8\]\[\left(- \infty, 7/8 \right) \cup \left(7/8, \infty \right)\]